# An Efficient Deep Learning Framework for Brain Stroke Diagnosis Using Computed Tomography (CT) Images

### v2

@@ 3.2 Image Preprocessing
Building a machine learning model requires data preprocessing since noise may be present in the images. Preprocessing techniques are essential for eliminating noisy data from the dataset and achieving high training accuracy. Preprocessing is necessary to increase efficiency since training a convolutional neural network directly with raw images may result in poor classification performance. A number of preprocessing techniques were used on the images before they were added to the model. This study included resizing and rescaling as pre-processing techniques. Convolutional Neural Networks require input images that have consistent sizes. Consequently, each image in the dataset was resized to have consistent dimensions of 224x224 pixels. Also we used data augmentation method. The technique of artificially creating additional data from preexisting data is known as data augmentation [32]. Through various adjustments like as rotation, flipping, scaling, and cropping, these techniques generate new images from the previously collected dataset. The model becomes less prone to overfitting and more resilient with the addition of new data, which finally improves the model’s performance on unobserved data.
@@ 3.3 Feature Extraction
Feature extraction involves identifying essential patterns or characteristics from raw data to make it suitable for machine learning models [33]. In deep learning, particularly with convolutional neural networks (CNNs), this process applies convolutional filters to input data to detect features like edges, textures, shapes, and high-level object structures. Each layer in the network captures increasingly abstract features, starting from simple patterns (e.g., lines and edges) in the initial layers to complex features (e.g., object parts or relationships) in deeper layers [34]. The process employs convolutions, pooling for dimensionality reduction, and activation functions to create a compact and meaningful representation of the input, enabling efficient data processing for classification or detection tasks [35].
@@ 3.3.1 MobileNetV2
MobileNetV2 is a lightweight deep neural network optimized for mobile and edge devices. It employs depthwise separable convolutions to reduce computational complexity and inverted residuals with linear bottlenecks to preserve representational power [36][37]. This architecture is designed for efficient feature extraction while maintaining high performance.
@@ 3.3.2 DensNet201
DenseNet201 is designed to connect each layer to every other layer in a feedforward fashion [38]. Unlike traditional architectures where layers connect only to subsequent layers, DenseNet enables feature reuse by concatenating feature maps from all preceding layers. This not only strengthens feature propagation but also reduces the number of parameters and mitigates vanishing gradients [39].
@@ 3.3.3 InceptionV3
InceptionV3 is a convolutional neural network model designed to process images at multiple scales simultaneously using inception modules [40]. These modules consist of parallel convolutional filters of different sizes (e.g., 1x1, 3x3, 5x5) that enable the model to capture both fine-grained and coarse features. It includes auxiliary classifiers and factorized convolutions to improve performance and reduce computational cost [41].
@@ 3.3.4 ResNet50
ResNet50 is a deep residual network with 50 layers known for introducing residual connections that bypass one or more layers [42]. These shortcuts allow gradients to flow directly through the network, solving the vanishing gradient problem and enabling the training of very deep models [43]. ResNet50 captures hierarchical spatial features effectively, making it suitable for image feature extraction [44].
@@ 3.3.5 Xception
Xception, short for ”Extreme Inception,” builds upon the Inception architecture by replacing standard convolutions with depthwise separable convolutions [45]. This decouples cross-channel and spatial correlations, significantly improving computational efficiency while maintaining model performance [46].
@@ 3.4 Feature Engineering
Feature engineering is the process of creating, transforming, or selecting relevant features from raw data to improve a model’s performance [47]. Feature selection is a subset of feature engineering that focuses on identifying the most relevant features for the model by eliminating redundant or irrelevant ones [48]. Feature reduction, on the other hand, reduces the dimensionality of the data while retaining essential information [49]. Together, these processes enhance model efficiency, reduce overfitting, and improve interpretability. In this research we employed BFO as feature selection technique and PCA and LDA as feature reduction techniques. Bacterial Foraging Optimization (BFO) is a bio-inspired optimization algorithm based on the foraging behavior of bacteria [50]. It is used for feature selection by finding the optimal subset of features that maximize classification accuracy or minimize a predefined cost function [51]. BFO mimics bacterial processes like chemotaxis (movement toward nutrients or away from harmful substances), reproduction, elimination, and dispersal [52]. Principal Component Analysis (PCA) is an unsupervised dimensionality reduction technique used to transform a high-dimensional feature space into a lower-dimensional space by identifying the most significant components (principal components) [53] [54]. These components are linear combinations of the original features and are selected to maximize variance while minimizing redundancy. PCA assumes that most of the information in the data lies in the directions of maximum variance. PCA is particularly useful for reducing noise and computational complexity in high-dimensional datasets [55]. Linear Discriminant Analysis (LDA) is a supervised feature reduction technique designed to maximize the separability between classes in the dataset [56]. Unlike PCA, which focuses on variance, LDA identifies the directions (linear discriminants) that maximize the ratio of between-class variance to within-class variance [57] [58]. This ensures that the projected features enhance class discrimination.
@@ 3.5 Classification Models
In this research we utilized a combination of pre-trained deep learning models for feature extraction and machine learning classifiers for categorizing stroke into distinct classes. After the features were extracted from CT images and applying feature optimization, leveraging the representation capabilities of pre-trained models, and subsequently fed into various machine learning classifiers to perform classification. We utilized seven Machine learning classifiers. The selected classifiers are diverse in their methodologies, offering unique approaches to classification. Support Vector Classifier (SVC) is a supervised machine learning algorithm that seeks to find the optimal hyperplane separating data points of different classes. By using kernel functions, SVC can handle both linear and non-linear classification tasks effectively, making it a versatile choice for classification problems [59, 60]. Random Forest (RF) is an ensemble learning method that constructs multiple decision trees during training and aggregates their predictions for final output. RF is particularly robust against overfitting, handles both categorical and numerical data well, and provides feature importance scores, offering insights into the most influential features [61]. Gaussian Naive Bayes (GNB) is based on Bayes theorem, this classifier assumes the independence of features within the dataset, which simplifies computation and makes it highly efficient for high-dimensional data. Despite its simplicity, GNB performs remarkably well in scenarios with normally distributed data [62]. Decision Tree (DT) is a tree-structured algorithm that splits the dataset into branches based on feature values, producing a highly interpretable model. However, decision trees are prone to overfitting, especially when the tree becomes overly complex, which necessitates pruning or ensemble methods for better generalization [63]. Extreme Gradient Boosting (XGBoost) is an advanced implementation of gradient boosting that improves classification performance by iteratively reducing errors from prior models. XGBoost is computationally efficient, handles missing data effectively, and supports regularization, making it a popular choice for structured datasets [64]. k-Nearest Neighbors (KNN) is a non-parametric algorithm that assigns class labels based on the majority vote of the nearest neighbors. While simple to implement, KNN can be computationally intensive for large datasets and is sensitive to the choice of distance metrics [65] [66]. Logistic Regression (LR) is a statistical model that estimates the probability of class membership using the logistic function. LR is effective for both binary and multi-class classification tasks, particularly when the relationship between features and the target variable is approximately linear [67]. Each of these classifiers, combined with the extracted features and optimized through the selected techniques, is evaluated for its ability to maximize accuracy and generalization.
@@ 4.3 Implementation Details
The proposed brain stroke diagnosis model begins with collecting and preparing datasets, Dataset-1 and Dataset-2, which are then manually curated to ensure data quality. This includes removing irrelevant and duplicate entries and incorporating additional data from local clinic to enhance the dataset’s diversity
Initialize: $DS_{1}$, $DS_{2}$, $PP$, $DA$, $CNN$, $FE$, $FEG$, $ML$, $EVAL$, $TC$, $SC$ ;
@@
[…]
@@
[…]
@@
[…]
Apply $PP$ on all images in $DS$ ;
@@
[…]
Apply $DA$ on $DS$ ;
@@
[…]
@@
[…]
       Compute $EVAL$ for $ML[k]$ ;
@@
Select the best model based on $EVAL$, $TC$, and $SC$ ;
Deinitialize: $DS_{1}$, $DS_{2}$, $PP$, $DA$, $CNN$, $FE$, $FEG$, $ML$, $EVAL$, $TC$, $SC$ ;
@@
and representativeness. The preprocessing stage involves resizing each image to consistent dimensions of 224×224 pixels for most CNN architectures and 299×299 pixels for Inception V3 and Xception. Rescaling is performed to normalize the pixel values to a range of [0, 1], ensuring compatibility with the input requirements of the candidate models. Following preprocessing, feature extraction is conducted using five pre-trained convolutional neural network (CNN) architectures: DenseNet201, InceptionV3, Xception, ResNet50, and MobileNet V2. The classification layers of these CNNs are removed, allowing the penultimate layer’s feature maps to be extracted. These extracted features are saved as high-dimensional vectors in a CSV format for further processing. The feature vectors are then subjected to either feature selection and dimensionality reduction techniques or directly used without feature selection. Feature reduction methods include Bacterial Foraging Optimization (BFO), A bio-inspired optimization technique for selecting the most relevant features. Principal Component Analysis (PCA) is a statistical method to reduce dimensionality while preserving variance. Linear Discriminant Analysis (LDA) is a technique to maximize class separability by projecting data into a lower-dimensional space. The processed features are subsequently fed into a set of traditional machine learning classifiers: Support Vector Classifier (SVC), Random Forest (RF), Gaussian Naive Bayes (GNB), Decision Tree (DT), Extreme Gradient Boosting (XGB), K-Nearest Neighbors (KNN), and Logistic Regression (LR). Each classifier is trained and evaluated to determine its performance in predicting stroke categories. Finally, the best-performing model is selected based on its performance metrics and is utilized to classify outcomes into three categories: hemorrhagic stroke, ischemic stroke, and normal condition. This end-to-end approach ensures a robust and accurate diagnosis system suitable for real-world clinical applications. The algorithm 1 shows the implementation procedure of the proposed brain stroke detection method.
@@ 4.4 Evaluation Matrices
[…]
@@ 5.1 Result Analysis
Fig. LABEL:fig:ac presents a comparison of classification accuracies across different machine learning models SVC, RF, GNB, DT, XGB, KNN, and LR integrated with MobileNetV2 under four feature selection techniques: no feature selection, BFO, PCA, and LDA. Each subfigure highlights the performance of these models under a specific feature selection approach. Fig. LABEL:fig:aa demonstrates the results with no feature selection, showing XGB achieving the highest accuracy (89.21%), while GNB performs the lowest. Fig. LABEL:fig:bb shows the results using the BFO method, where KNN achieves the highest accuracy (90.66%). Fig. LABEL:fig:cc uses PCA, with SVC leading in performance (86.03%). Fig. LABEL:fig:dd incorporates LDA, where SVC achieves the best accuracy (97.93%). These results suggest that combining MobileNetV2 with LDA and SVC yields the most accurate classification among the tested setups.
@@
Fig. LABEL:fig:lc displays learning curves that compare the performance of different feature optimization techniques applied to MobileNetV2 in terms of accuracy scores across varying training set sizes. Among the approaches, Fig. LABEL:fig:hh LDA + SVC outperforms the others, demonstrating the highest and most stable cross-validation accuracy. In contrast, Fig. LABEL:fig:gg PCA + SVC shows moderate performance but suffers from higher variance. Fig. LABEL:fig:ff BFO + KNN and Fig. LABEL:fig:ee XGB without optimization exhibit relatively lower accuracy and greater fluctuations. While all models improve with increased training data, LDA + SVC proves to be the most effective combination, offering superior and consistent classification performance.
@@
Fig. LABEL:fig:cm illustrates confusion matrices comparing different feature optimization techniques applied to MobileNetV2 for classifying Normal, Ischemic, and Hemorrhagic cases. Among the methods, Support Vector Classifier (SVC) with Linear Discriminant Analysis (LDA) outperforms the others, achieving the highest accuracy with 98% for the Ischemic class and 95% for the Hemorrhagic class. Other techniques, such as Bacterial Feature Optimization (BFO) + K-Nearest Neighbors (KNN) and Principal Component Analysis (PCA) + SVC, show relatively lower performance, with misclassification rates between 11% and 30% in certain cases. The no-feature-selection approach with SVC demonstrates moderate classification accuracy, but LDA + SVC proves to be the best-performing combination.
@@
Fig. LABEL:fig:rc presents ROC curves comparing different feature optimization techniques applied to MobileNetV2 with various classifiers. Among them, Fig. LABEL:fig:pp, the LDA + SVC combination, demonstrates the best overall performance, achieving the highest AUC values across all three classes. Compared to Fig. LABEL:fig:oo, the PCA + SVC combination, and Fig. LABEL:fig:nn, the BFO + KNN approach, the LDA-based method exhibits superior class separability, particularly for class 1 and class 2, which are typically more challenging to classify. On the other hand, Fig. LABEL:fig:mm, XGBoost without optimization, struggles with lower AUC scores, especially for class 2 and 3. These results indicate that LDA effectively enhances feature discrimination, making its combination with SVC the most reliable approach for our brain stroke classification task.
@@
Table 9 compares the computational complexity of various deep learning techniques for stroke detection, analyzing both time and space requirements. Among the models, InceptionV3+SVC is the fastest, with a time complexity of 132ms ± 10.2ms, while Xception+KNN demand the most memory ( 1962.64 MiB PM). DenseNet201+SVC has the lowest peak memory usage (1493.54 MiB) but suffers from a higher inference time (309ms ± 128ms). MobileNetV2+SVC achieves a strong balance with 186ms ± 48.3ms time complexity and a moderate memory footprint (PM: 1546.84 MiB, INC: 569.09 MiB).
@@
Beyond computational efficiency, MobileNetV2+SVC stands out as the best model in terms of overall performance, excelling in accuracy and other classification metrics while maintaining a good trade-off between speed and memory usage. It’s relatively low computational cost makes it a practical choice for real-time stroke detection applications.
5.2 Discussion and Limitations
In this study, we proposed a feature extraction-based approach utilizing pre-trained deep learning models combined with various classifiers and feature optimization techniques to enhance classification performance.
@@ 5.2 Discussion and Limitations
A. Abumihsan et al. [20] 2000 & 2501 2 99.21% Moderate
M.M Hossain et al. [21] 770 & 2501 2 96.61% High
I. Tahyudin et al. [22] 2501 2 95.00% High
UmaMaheswaran et al. [23] 2501 2 97.00% Low
M.A. Saleem et al. [24] 2501 2 96.50% High
M. Sabir et al. [25] 10000 2 96.50% High
S. Prasher et al. [26] 2501 2 98.72% High
C.D. Kulathilake et al. [27] 4767 2 98.02% High
R. Raj et al. [28] 2431 2 92.00% High
A. Diker et al. [29] 2501 2 97.06% High
@@
Instead of relying on computationally expensive end-to-end deep learning models, we leveraged transfer learning to extract meaningful features from brain imaging data, which were then fed into traditional machine learning classifiers for stroke classification. Our extensive experimentation with multiple model-classifier combinations revealed that MobileNetV2+LDA+SVC was the best-performing model. This combination achieved the highest accuracy, precision, recall, and F1-score while maintaining a strong trade-off between computational efficiency and classification performance. Additionally, the model demonstrated robustness across different feature selection techniques, further validating its reliability for stroke detection. Compared to other deep learning models, MobileNetV2 is a lightweight architecture, making it a practical choice for real-world applications where computational resources may be limited.
We compared our findings with prior research on brain stroke detection using computed tomography (CT) images; our model outperformed most existing studies in this domain. While some studies, such as [20, 26, 27] used Hybrid CBAM, EfficientNet-B0, and DenseNet201, reported slightly higher accuracy. it is important to note that these models likely employed end-to-end training strategies that require extensive computational resources and large-scale datasets for optimal performance. In contrast, our feature extraction-based method offers a more efficient and scalable solution, as it eliminates the need for training deep models from scratch. Additionally, many of the high-performing models in prior studies do not report computational complexity, making it difficult to assess their feasibility for deployment in real-time clinical settings. By incorporating feature optimization techniques such as PCA, LDA, and BFO, our approach not only enhances classification accuracy but also reduces the risk of overfitting and improves generalization across different dataset sizes. Another key advantage of our model is its ability to balance high classification performance with computational efficiency, making it a viable option for deployment in medical imaging systems and resource-constrained environments.

### v3

@@ 3.2 Image Preprocessing
Building a machine learning model requires data preprocessing since noise may be present in the images. Preprocessing techniques are essential for eliminating noisy data from the dataset and achieving high training accuracy. Preprocessing is necessary to increase efficiency since training a convolutional neural network directly with raw images may result in poor classification performance. A number of preprocessing techniques were used on the images before they were added to the model. This study included resizing and rescaling as pre-processing techniques. Convolutional Neural Networks require input images that have consistent sizes. Consequently, each image in the dataset was resized to have consistent dimensions of 224x224 pixels. Also we used data augmentation method. The technique of artificially creating additional data from preexisting data is known as data augmentation [strawberry]. Through various adjustments like as rotation, flipping, scaling, and cropping, these techniques generate new images from the previously collected dataset. The model becomes less prone to overfitting and more resilient with the addition of new data, which finally improves the model’s performance on unobserved data.
@@ 3.3 Feature Extraction
Feature extraction involves identifying essential patterns or characteristics from raw data to make it suitable for machine learning models [featureextraction3]. In deep learning, particularly with convolutional neural networks (CNNs), this process applies convolutional filters to input data to detect features like edges, textures, shapes, and high-level object structures. Each layer in the network captures increasingly abstract features, starting from simple patterns (e.g., lines and edges) in the initial layers to complex features (e.g., object parts or relationships) in deeper layers [featureextraction1]. The process employs convolutions, pooling for dimensionality reduction, and activation functions to create a compact and meaningful representation of the input, enabling efficient data processing for classification or detection tasks [featureextraction2].
@@ 3.3.1 MobileNetV2
MobileNetV2 is a lightweight deep neural network optimized for mobile and edge devices. It employs depthwise separable convolutions to reduce computational complexity and inverted residuals with linear bottlenecks to preserve representational power [mnt1][mnt2]. This architecture is designed for efficient feature extraction while maintaining high performance.
@@ 3.3.2 DensNet201
DenseNet201 is designed to connect each layer to every other layer in a feedforward fashion [dense1]. Unlike traditional architectures where layers connect only to subsequent layers, DenseNet enables feature reuse by concatenating feature maps from all preceding layers. This not only strengthens feature propagation but also reduces the number of parameters and mitigates vanishing gradients [dense2].
@@ 3.3.3 InceptionV3
InceptionV3 is a convolutional neural network model designed to process images at multiple scales simultaneously using inception modules [inception1]. These modules consist of parallel convolutional filters of different sizes (e.g., 1x1, 3x3, 5x5) that enable the model to capture both fine-grained and coarse features. It includes auxiliary classifiers and factorized convolutions to improve performance and reduce computational cost [inception2].
@@ 3.3.4 ResNet50
ResNet50 is a deep residual network with 50 layers known for introducing residual connections that bypass one or more layers [rn1]. These shortcuts allow gradients to flow directly through the network, solving the vanishing gradient problem and enabling the training of very deep models [rn2]. ResNet50 captures hierarchical spatial features effectively, making it suitable for image feature extraction [rn3].
@@ 3.3.5 Xception
Xception, short for ”Extreme Inception,” builds upon the Inception architecture by replacing standard convolutions with depthwise separable convolutions [xception1]. This decouples cross-channel and spatial correlations, significantly improving computational efficiency while maintaining model performance [xception2].
@@ 3.4 Feature Engineering
Feature engineering is the process of creating, transforming, or selecting relevant features from raw data to improve a model’s performance [featureengineering]. Feature selection is a subset of feature engineering that focuses on identifying the most relevant features for the model by eliminating redundant or irrelevant ones [featureseleciton]. Feature reduction, on the other hand, reduces the dimensionality of the data while retaining essential information [featurereduction]. Together, these processes enhance model efficiency, reduce overfitting, and improve interpretability. In this research we employed BFO as feature selection technique and PCA and LDA as feature reduction techniques. Bacterial Foraging Optimization (BFO) is a bio-inspired optimization algorithm based on the foraging behavior of bacteria [bfo1]. It is used for feature selection by finding the optimal subset of features that maximize classification accuracy or minimize a predefined cost function [bfo2]. BFO mimics bacterial processes like chemotaxis (movement toward nutrients or away from harmful substances), reproduction, elimination, and dispersal [bfo3]. Principal Component Analysis (PCA) is an unsupervised dimensionality reduction technique used to transform a high-dimensional feature space into a lower-dimensional space by identifying the most significant components (principal components) [pca1] [pca2]. These components are linear combinations of the original features and are selected to maximize variance while minimizing redundancy. PCA assumes that most of the information in the data lies in the directions of maximum variance. PCA is particularly useful for reducing noise and computational complexity in high-dimensional datasets [Sentiment]. Linear Discriminant Analysis (LDA) is a supervised feature reduction technique designed to maximize the separability between classes in the dataset [lda1]. Unlike PCA, which focuses on variance, LDA identifies the directions (linear discriminants) that maximize the ratio of between-class variance to within-class variance [lda2] [lda3]. This ensures that the projected features enhance class discrimination.
@@ 3.5 Classification Models
In this research we utilized a combination of pre-trained deep learning models for feature extraction and machine learning classifiers for categorizing stroke into distinct classes. After the features were extracted from CT images and applying feature optimization, leveraging the representation capabilities of pre-trained models, and subsequently fed into various machine learning classifiers to perform classification. We utilized seven Machine learning classifiers. The selected classifiers are diverse in their methodologies, offering unique approaches to classification. Support Vector Classifier (SVC) is a supervised machine learning algorithm that seeks to find the optimal hyperplane separating data points of different classes. By using kernel functions, SVC can handle both linear and non-linear classification tasks effectively, making it a versatile choice for classification problems [svm1, svm2]. Random Forest (RF) is an ensemble learning method that constructs multiple decision trees during training and aggregates their predictions for final output. RF is particularly robust against overfitting, handles both categorical and numerical data well, and provides feature importance scores, offering insights into the most influential features [rf1]. Gaussian Naive Bayes (GNB) is based on Bayes theorem, this classifier assumes the independence of features within the dataset, which simplifies computation and makes it highly efficient for high-dimensional data. Despite its simplicity, GNB performs remarkably well in scenarios with normally distributed data [nb1]. Decision Tree (DT) is a tree-structured algorithm that splits the dataset into branches based on feature values, producing a highly interpretable model. However, decision trees are prone to overfitting, especially when the tree becomes overly complex, which necessitates pruning or ensemble methods for better generalization [dt1]. Extreme Gradient Boosting (XGBoost) is an advanced implementation of gradient boosting that improves classification performance by iteratively reducing errors from prior models. XGBoost is computationally efficient, handles missing data effectively, and supports regularization, making it a popular choice for structured datasets [xgb1]. k-Nearest Neighbors (KNN) is a non-parametric algorithm that assigns class labels based on the majority vote of the nearest neighbors. While simple to implement, KNN can be computationally intensive for large datasets and is sensitive to the choice of distance metrics [knn1] [maternal]. Logistic Regression (LR) is a statistical model that estimates the probability of class membership using the logistic function. LR is effective for both binary and multi-class classification tasks, particularly when the relationship between features and the target variable is approximately linear [lr1]. Each of these classifiers, combined with the extracted features and optimized through the selected techniques, is evaluated for its ability to maximize accuracy and generalization.
@@ 4.3 Implementation Details
The proposed brain stroke diagnosis model begins with collecting and preparing datasets, Dataset-1 and Dataset-2, which are then manually curated to ensure data quality. This includes removing irrelevant and duplicate entries and incorporating additional data from the local clinic to enhance the dataset’s diversity and representativeness. The preprocessing stage involves resizing each image to consistent dimensions of 224×224 pixels for most CNN architectures and 299×299 pixels for Inception V3 and Xception. Rescaling is performed to normalize the pixel values to a range of [0, 1], ensuring compatibility with the input requirements of the candidate models. Following preprocessing, feature extraction is conducted using five pre-trained convolutional neural network (CNN) architectures: DenseNet201, InceptionV3, Xception, ResNet50, and MobileNet V2. The classification layers of these CNNs are removed, allowing the penultimate layer’s feature maps to be extracted. These extracted features are saved as high-dimensional vectors in a CSV format for further processing. The feature vectors are then subjected to either feature selection and dimensionality reduction techniques or directly used without feature selection. Feature reduction methods include Bacterial Foraging Optimization (BFO), A bio-inspired optimization technique for selecting the most relevant features. Principal Component Analysis (PCA) is a statistical method to reduce dimensionality while preserving variance. Linear Discriminant Analysis (LDA) is a technique to maximize class separability by projecting data into a lower-dimensional space. The processed features are subsequently fed into a set of traditional machine learning classifiers: Support Vector Classifier (SVC), Random Forest (RF), Gaussian Naive Bayes (GNB), Decision Tree (DT), Extreme Gradient Boosting (XGB), K-Nearest Neighbors (KNN), and Logistic Regression (LR). Each classifier is trained and evaluated to determine its performance in predicting stroke categories. Finally, the best-performing model is selected based on its performance metrics and is utilized to classify outcomes into three categories: hemorrhagic stroke, ischemic stroke, and normal condition. This end-to-end approach ensures a robust and accurate diagnosis system suitable for real-world clinical applications. The algorithm 1 shows the implementation procedure of the proposed brain stroke detection method.
Initialize: $DS_{1}$, $DS_{2}$, PPPP, DADA, CNNCNN, FEFE, FEGFEG, MLML, EVALEVAL, TCTC, SCSC ;
@@
$DS_{2}\leftarrow\text{Brain Stroke Prediction CT Scan Image Dataset}$ ;
@@
$DS^{\prime}\leftarrow DS_{1}[\text{Selected Normal}]\cup DS_{2}[\text{Ischemic, Hemorrhagic}]$ ;
@@
$PP\leftarrow[\text{Resizing, Rescaling}]$ ;
Apply PPPP on all images in DSDS ;
@@
$DA\leftarrow[\text{Random Flipping, Rotation, Scaling, and Cropping}]$ ;
Apply DADA on DSDS ;
@@
$CNN\leftarrow[\text{MobileNetV2, ResNet50, DenseNet201, InceptionV3, Xception}]$ ;
@@
  $EVAL\leftarrow[\text{Accuracy, Precision, Recall, F1-Score, AUC}]$ ;
  Compute EVALEVAL for $ML[k]$ ;
@@
Select the best model based on EVALEVAL, TCTC, and SCSC ;
Deinitialize: $DS_{1}$, $DS_{2}$, PPPP, DADA, CNNCNN, FEFE, FEGFEG, MLML, EVALEVAL, TCTC, SCSC ;
@@ 4.4 Evaluation Matrices
    $\scriptsize\text{F1 Score}=2\times\frac{\text{Recall}\times\text{Precision}}{\text{Recall}+\text{Precision}}\times 100$ (4)
@@ 5.1 Result Analysis
(a) Bar Charts compare the classification performance of different feature selection techniques applied with MobileNetV2 and different classifiers: (a) No Feature Selection, (b) Bacterial Foraging Optimization (BFO), (c) Principal Component Analysis (PCA), and (d) Linear Discriminant Analysis (LDA). Each feature selection technique shows the comparison of accuracy across seven different classifiers.
(b) Learning curves comparing the performance of different feature optimization techniques employed with MobileNetV2 in terms of accuracy score with varying training set sizes. (a) Extreme Gradient Boosting with No Feature Optimization, (b) Bacterial Foraging Optimization with k-Nearest Neighbors, (c) Principal Component Analysis with Support Vector Classifier, (d) Linear Discriminant Analysis and Support Vector Classifier.
@@
Fig. 23(a) presents a comparison of classification accuracies across different machine learning models SVC, RF, GNB, DT, XGB, KNN, and LR integrated with MobileNetV2 under four feature selection techniques: no feature selection, BFO, PCA, and LDA. Each subfigure highlights the performance of these models under a specific feature selection approach. Fig. LABEL:fig:aa demonstrates the results with no feature selection, showing XGB achieving the highest accuracy (89.21%), while GNB performs the lowest. Fig. LABEL:fig:bb shows the results using the BFO method, where KNN achieves the highest accuracy (90.66%). Fig. LABEL:fig:cc uses PCA, with SVC leading in performance (86.03%). Fig. LABEL:fig:dd incorporates LDA, where SVC achieves the best accuracy (97.93%). These results suggest that combining MobileNetV2 with LDA and SVC yields the most accurate classification among the tested setups.
(a) Confusion matrices illustrating the performance of different feature selection techniques applied with MobileNetV2 and different classifiers: (a) No feature selection with Extreme Gradient Boosting, (b) Bacterial Feature Optimization with K-Nearest Neighbors, (c) Principal Component Analysis with Support Vector Classifier, and (d) Linear Discriminant Analysis with Support Vector Classifier. Each matrix shows the classification performance across three classes: Normal, Ischemic, and Hemorrhagic, with values representing the number of correctly and incorrectly classified instances.
(b) ROC Curve illustrating the performance of different feature optimization techniques applied with MobileNetV2 and various classifiers: (a) No feature selection with SVC, (b) Bacterial Foraging Optimization with K-Nearest Neighbors, (c) Principal Component Analysis with SVC, and (d) Linear Discriminant Analysis with Support Vector Classifier. Figures show the ROC Curve across three classes: 0 (Normal), 1 (Ischemic), 2 (Hemorrhagic), with values.
@@
Fig. 23(b) displays learning curves that compare the performance of different feature optimization techniques applied to MobileNetV2 in terms of accuracy scores across varying training set sizes. Among the approaches, Fig. LABEL:fig:hh LDA + SVC outperforms the others, demonstrating the highest and most stable cross-validation accuracy. In contrast, Fig. LABEL:fig:gg PCA + SVC shows moderate performance but suffers from higher variance. Fig. LABEL:fig:ff BFO + KNN and Fig. LABEL:fig:ee XGB without optimization exhibit relatively lower accuracy and greater fluctuations. While all models improve with increased training data, LDA + SVC proves to be the most effective combination, offering superior and consistent classification performance.
@@
Fig. 27(a) illustrates confusion matrices comparing different feature optimization techniques applied to MobileNetV2 for classifying Normal, Ischemic, and Hemorrhagic cases. Among the methods, Support Vector Classifier (SVC) with Linear Discriminant Analysis (LDA) outperforms the others, achieving the highest accuracy with 98% for the Ischemic class and 95% for the Hemorrhagic class. Other techniques, such as Bacterial Feature Optimization (BFO) + K-Nearest Neighbors (KNN) and Principal Component Analysis (PCA) + SVC, show relatively lower performance, with misclassification rates between 11% and 30% in certain cases. The no-feature-selection approach with SVC demonstrates moderate classification accuracy, but LDA + SVC proves to be the best-performing combination.
@@
Fig. 27(b) presents ROC curves comparing different feature optimization techniques applied to MobileNetV2 with various classifiers. Among them, Fig. LABEL:fig:pp, the LDA + SVC combination, demonstrates the best overall performance, achieving the highest AUC values across all three classes. Compared to Fig. LABEL:fig:oo, the PCA + SVC combination, and Fig. LABEL:fig:nn, the BFO + KNN approach, the LDA-based method exhibits superior class separability, particularly for class 1 and class 2, which are typically more challenging to classify. On the other hand, Fig. LABEL:fig:mm, XGBoost without optimization, struggles with lower AUC scores, especially for class 2 and 3. These results indicate that LDA effectively enhances feature discrimination, making its combination with SVC the most reliable approach for our brain stroke classification task.
@@
Table 9 compares the computational complexity of various deep learning techniques for stroke detection, analyzing both time and space requirements. Among the models, InceptionV3+SVC is the fastest, with a time complexity of 132ms ± 10.2ms, while Xception+KNN demands the most memory ( 1962.64 MiB PM). DenseNet201+SVC has the lowest peak memory usage (1493.54 MiB) but suffers from a higher inference time (309ms ± 128ms). MobileNetV2+SVC achieves a strong balance with 186ms ± 48.3ms time complexity and a moderate memory footprint (PM: 1546.84 MiB, INC: 569.09 MiB). Beyond computational efficiency, MobileNetV2+SVC stands out as the best model in terms of overall performance, excelling in accuracy and other classification metrics
@@
A. Abumihsan et al. [abumihsan] 2000 & 2501 2 99.21% Moderate
M.M Hossain et al. [hossainmaruf] 770 & 2501 2 96.61% High
I. Tahyudin et al. [tahyudin2025high] 2501 2 95.00% High
UmaMaheswaran et al. [umamaheswaran2024enhanced] 2501 2 97.00% Low
M.A. Saleem et al. [saleem] 2501 2 96.50% High
M. Sabir et al. [sabir] 10000 2 96.50% High
S. Prasher et al. [prasher2024brain] 2501 2 98.72% High
C.D. Kulathilake et al. [kulathilake] 4767 2 98.02% High
R. Raj et al. [raj2023strokevit] 2431 2 92.00% High
A. Diker et al. [diker] 2501 2 97.06% High
@@
while maintaining a good trade-off between speed and memory usage. Its relatively low computational cost makes it a practical choice for real-time stroke detection applications.
5.2 Discussion and Limitations
In this study, we proposed a feature extraction-based approach utilizing pre-trained deep learning models combined with various classifiers and feature optimization techniques to enhance classification performance. Instead of relying on computationally expensive end-to-end deep learning models, we leveraged transfer learning to extract meaningful features from brain imaging data, which were then fed into traditional machine learning classifiers for stroke classification. Our extensive experimentation with multiple model-classifier combinations revealed that MobileNetV2+LDA+SVC was the best-performing model. This combination achieved the highest accuracy, precision, recall, and F1-score while maintaining a strong trade-off between computational efficiency and classification performance. Additionally, the model demonstrated robustness across different feature selection techniques, further validating its reliability for stroke detection. Compared to other deep learning models, MobileNetV2 is a lightweight architecture, making it a practical choice for real-world applications where computational resources may be limited.
We compared our findings with prior research on brain stroke detection using computed tomography (CT) images; our model outperformed most existing studies in this domain. While some studies, such as [abumihsan, prasher2024brain, kulathilake] used Hybrid CBAM, EfficientNet-B0, and DenseNet201, reported slightly higher accuracy. it is important to note that these models likely employed end-to-end training strategies that require extensive computational resources and large-scale datasets for optimal performance. In contrast, our feature extraction-based method offers a more efficient and scalable solution, as it eliminates the need for training deep models from scratch. Additionally, many of the high-performing models in prior studies do not report computational complexity, making it difficult to assess their feasibility for deployment in real-time clinical settings. By incorporating feature optimization techniques such as PCA, LDA, and BFO, our approach not only enhances classification accuracy but also reduces the risk of overfitting and improves generalization across different dataset sizes. Another key advantage of our model is its ability to balance high classification performance with computational efficiency, making it a viable option for deployment in medical imaging systems and resource-constrained environments.